OpenStudy (nathalyn):

5x + 10x = 75 Please help!

Mehek (mehek14):

\(5x +10x = 5+10\)

OpenStudy (nathalyn):

I need the whole process

Mehek (mehek14):

first add 5x + 10x

Mehek (mehek14):

which is the same as 5 + 10 but add the x

OpenStudy (anonymous):

5x+10x=75 15x=75 x=5

OpenStudy (mrnood):

you don't need the whole process GIVEN to you you need to DO the whole process so YOU understand it

OpenStudy (mrnood):

you don't need the whole process GIVEN to you you need to DO the whole process so YOU understand it

Mehek (mehek14):

are you going to answer 5x +10x?

OpenStudy (mrnood):

probably not now because @paolaxx33 gave him the answer - which was all he really wanted

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NathalyN I need the whole process \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (nathalyn):

um im not a he I'm a her!!! @MrNood

OpenStudy (mrnood):

ok - sorry but since you have the answer can you solve THIS for yourself? it is EXACTLY the same process... 12x +8x = 60 I f so you have learnt something . If not you are wasting your time here

OpenStudy (nathalyn):

you know @MrNood I haev been wacthing you and I don't want to very mean I dont much when you to the people dont be rude ! okay people dont likt and i dont like and no I wil not @MrNood :/

OpenStudy (mrnood):

If you spend your school life copying other people's answers you are wasting the best opportunity you will ever have. Try to understand that. And try to make some contribution to your own questions - rather than requesting someone else to give you the 'process' you can copy into your homework/test/whatever.

OpenStudy (anonymous):

add apples with apples and oranges to oranges. 5x + 10x is like 5+!0 which =15 or with the x. 15x. so now 15x=75. divide both sides by fifteen so you only have x on the one side. so 75/15 = 5. your answer is x=5

OpenStudy (anonymous):

5x+10x=75 Simplify 5x+10x=75 (5x+10x)=75(Combine Like Terms) 15x=75 Divide both sides by 15. 15x/15 = 75 /15 x=5